# Multi-level Temporal-channel Speaker Retrieval for Zero-shot Voice Conversion

### v2

@@ Abstract
Zero-shot voice conversion (VC) converts source speech into the voice of any desired speaker using only one utterance of the speaker without requiring additional model updates. Typical methods use a speaker representation from a pre-trained speaker verification (SV) model or learn speaker representation during VC training to achieve zero-shot VC. However, existing speaker modeling methods overlook the variation of speaker information richness in temporal and frequency channel dimensions of speech. This insufficient speaker modeling hampers the ability of the VC model to accurately represent unseen speakers who are not in the training dataset. In this study, we present a robust zero-shot VC model with multi-level temporal-channel retrieval, referred to as MTCR-VC. Specifically, to flexibly adapt to the dynamic-variant speaker characteristic in the temporal and channel axis of the speech, we propose a novel fine-grained speaker modeling method, called temporal-channel retrieval (TCR), to find out when and where speaker information appears in speech. It retrieves variable-length speaker representation from both temporal and channel dimensions under the guidance of a pre-trained SV model. Besides, inspired by the hierarchical process of human speech production, the MTCR speaker module stacks several TCR blocks to extract speaker representations from multi-granularity levels. Furthermore, to achieve better speech disentanglement and reconstruction, we introduce a cycle-based training strategy to simulate zero-shot inference recurrently. We adopt perpetual constraints on three aspects, including content, style, and speaker, to drive this process. Experiments demonstrate that MTCR-VC is superior to the previous zero-shot VC methods in modeling speaker timbre while maintaining good speech naturalness.
@@ I Introduction
Voice conversion (VC) aims to convert the source speech to a target speaker without changing the linguistic content. VC has been deployed to many applications, including dubbing, live broadcast, voice anonymization, and pronunciation correction. In recent years, neural networks-based VC systems, such as generative adversarial network (GAN) [1], variational auto-encoder (VAE) [2], and recognition-synthesis autoencoder framework [3] are widely adopted. But these VC models can only convert source speech to that of predefined target speakers. Meanwhile, these VC models usually require a large number of speech recordings of the target speaker for model training. As data collection is expensive and time-consuming, building a high-quality VC system with minimal data requirements is more practical for real-world applications. Consequently, zero-shot VC, also called any-to-any VC, which aims to convert source speech to that of any speaker given only one utterance of the speaker, has drawn much attention recently. In this paper, we focus on the problem of zero-shot VC.
@@
As only one utterance is available, one of the key challenges in zero-shot VC is how to effectively capture the target speaker’s timbre to achieve good speaker similarity in the converted speech. Using a look-up table (LuT) to represent speaker identity [4, 5, 6] is a common practice but is limited to predefined speakers in the training data. Many recent studies attempt to directly capture speaker timbre from the limited speech of the target speaker. An intuitive approach is to leverage a pre-trained robust speaker verification (SV) model to extract utterance-level speaker representation [7, 8, 9, 10]. However, the SV model trained on a large number of speakers and recording conditions is optimized for speaker classification but not for the perceptual speaker similarity in the sense of human hearing. Instead of using an external SV model, many studies [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21] decompose speech into linguistic content and speaker timbre, and even speaking style, aiming to separate utterance-level speaker timbre from other speech components. In the above approaches, speaker timbre is considered static and time-independent and modeled as a single coarse-grained fixed-length vector. To capture the dynamically varying speaker characteristics within an utterance, some recent studies [22, 23, 24, 25, 26] focus on modeling more fine-grained speaker representation, which extracts speaker timbre from multiple aspects, covering multi-level and time-varying speaker representation. Based on the U-net [27] structure, Li et al. [28], Wu et al. [22], and Li et al. [26] extract utterance-level speaker representations from multiple stacking layers and feed them to corresponding decoder layers. To access time-varying speaker information, some studies [25, 23, 24] extract variable-length speaker representation and fuse it into the converted speech according to the content-based alignment between source speech and target speaker speech.
@@
While the above progress has been made, existing methods of modeling fine-grained target speaker timbre are insufficient to capture the dynamic variation characteristic of speaker timbre. The above speaker representation is usually extracted from the time-frequency space of speech, where a speech spectrogram includes both temporal and frequency channel dimensions. There are differences in speaker timbre richness in both temporal and channel dimensions of speech. Insufficient speaker modeling in these two dimensions may cause the unstable performance of the VC model to represent unseen speakers. Moreover, speech production research [29] shows that the frequency distributions of speech from different speakers lead to varying speaker timbre information in frequency channels [30, 31, 32]. Speech content, such as vowels, consonants, and para-linguistic features, carry distinct speaker timbre information reflected in temporal and frequency channel dimensions while silent speech segments apparently convey no speaker timbre information [25]. On the other hand, the human speech production mechanism is hierarchical [33, 29] in nature from long-term airflow generation to fine-grained phoneme-related articulator movements and vocal filtering. Speaker-related information thus varies at different stages of speech production with different temporal and channel granularities. Finally, different speech factors such as linguistic content, speaking style, and speaker timbre are highly entangled in speech, and factor disentanglement is necessary for better speaker timbre extraction. Compared with utterance-level speaker modeling, time-varying fine-grained speaker modeling is able to transfer more speaker information from the target speaker’s speech to the converted speech. But this exacerbates the difficulty of disentanglement [25, 23].
@@
To address the aforementioned problems in zero-shot VC, we propose a novel fine-grained speaker modeling method called temporal-channel retrieval (TCR), which captures the dynamic variation of speaker timbre in both temporal and channel dimensions. Specifically, with the help of the attention mechanism, the speaker embedding from the pre-trained SV model is used as a query to retrieve speaker timbre information in variable-length speaker representation from the target speaker. Inspired by the hierarchical nature of speech production, we employ multi-level TCR (MTCR) in an encoder-decoder based U-net structure for voice conversion [27] where fine-grained speaker information in different granularities is retrieved from different encoder layers. In the decoder, these multi-level speaker representations are then fused with content and style representations from the source speech to generate target speech. Furthermore, in order to perform better speech disentanglement, we introduce a cycle-based training strategy to simulate zero-shot inference in a recurrent fashion. Perpetual constraints [34] on three aspects, including content, style, and speaker, are adopted to drive this process. Zero-shot voice conversion experiments show that the proposed MTCR-VC approach generalizes well to cross-set speakers with superior speaker similarity. Audio samples can be found on our demo page. [^1] https://kerwinchao.github.io/demo_zslvc/
@@
  • We propose a novel multi-level framework MTCR-VC for fine-grained speaker modeling in zero-shot VC. The multi-level temporal-channel retrieval is designed to extract fine-grained speaker representations under the query of SV’s speaker embedding in both temporal and channel dimensions and hierarchically capture speaker timbre from different granularities.
@@ II Related work
This section will review related works on voice conversion and the typical zero-shot VC frameworks. Besides, we will also review the utilization of the attention mechanism in VC.
@@ II-A Voice Conversion
VC approaches can be categorized into one-to-one, many-to-many, or any-to-any VC based on the number of speakers that the VC model supports. Conventional VC approaches focus on one-to-one VC, which align acoustic features of parallel data between a pair of source-target speakers and perform frame-wise mapping using Gaussian mixture models [35, 36] and neural networks [37, 38]. With the high cost of parallel data, recent studies mainly use non-parallel data for voice conversion. GAN [4, 5] and VAE [1] are further proposed to support many-to-many VC. With the help of text supervision, phonetic posteriorgram (PPG) and neural bottleneck feature (BNF) [3] computed from an automatic speech recognition (ASR) model are assumed to be speaker-independent content representations. And thus PPG- and BN-based recognition-synthesis framework is widely used in any-to-many VC. Besides, many disentanglement methods based on information bottleneck [7], perturbation [8, 16], and representation constraints [17, 13] decompose speech into content, speaker timbre, and speaking style to achieve any-to-any VC, also called zero-shot VC. In this work, based on the recognition-synthesis framework, we propose a new method for zero-shot VC.
@@ II-B Typical Zero-shot VC Framework
The key to voice conversion is to decompose speech into linguistic content and speaker timbre, or even speaking style components. Ideally, only the content and style components from the source speech are transferred to the target speech with the target speaker’s timbre. In many studies [11, 12, 23, 10], style is not explicitly considered in the disentanglement. In zero-shot scenarios, the speaker timbre representation is extracted from a short speech utterance of the target speaker, and then the speaker representation is fused with the extracted content and style of the source speech to generate the target speech. The extraction of speaker timbre can be done at utterance level or fine-grained level, as shown in Fig. 1.
@@ II-B2 Fine-grained Speaker Modeling
Fine-grained speaker modeling differs from utterance-level speaker modeling in that it accounts for changes in speaker timbre over time and considers dynamic variation characteristics. Fig. 1(b) and (c) illustrate two categories of fine-grained speaker modeling: time-varying and multi-level. In the time-varying speaker modeling [25, 23, 24] as shown in Fig.1(b), variable-length speaker representation is extracted from the target speaker’s speech and then fused with the content and style of the source speech based on the alignment between them. On the other hand, multi-level speaker modeling [28, 26] shown in Fig.1(c) typically adopts a U-net structure to extract multiple utterance-level speaker representations from stacked layers and then feeds them to the corresponding decoder layers. Recent studies [23, 25, 24] attempt to employ time-varying and multi-level speaker modeling simultaneously.
Despite this progress, existing methods are insufficient to capture the dynamic variation characteristic of the speaker timbre, which varies across not only temporal but also channel regions of speech at different granularities. This insufficient may cause unstable speaker modeling and result in low speaker similarity. Besides, in a time-varying fine-grained speaker modeling framework, speech factors are much easier to entangle with each other, making achieving zero-shot more challenging. In this paper, we focus on fine-grained speaker modeling in our proposed zero-shot VC framework, which models speaker timbre from both temporal and channel dimensions at different granularities. Comprehensive representations of speech components and a well-designed training strategy ensure the decoupling capabilities of our model.
@@ II-C Attention Mechanism in Speaker Modeling
In recent years, many attention mechanisms have been proposed, such as spatial attention [39], channel attention [40], temporal attention [41], and their combination [42, 43]. Some attention mechanisms have also been explored in speaker modeling. In Desplanques et al. [32], Squeeze-and-excitation [40] block (SE) is introduced to aggregate speaker information along the channel axis. Based on Desplanques et al. [32], Liu et al. [30] introduces multi-scale channel attention to better capture the speaker characteristics at any local frequency region. Besides, Sang et al.[31] improve the process of extracting utterance-level speaker representation from frame-level speaker representation by modified channel attention driven by a discrete cosine transform. For zero-shot VC, following Desplanques et al. [32], Du et al. [21] and Choi et al. [16] utilize SE block to improve the speaker modeling ability for zero-shot VC. CA-VC uses channel attention to improve content learning and encourage utterance-level speaker modeling. Besides, Retriever [24] uses cross-attention [44] learn a set of permutation invariant tokens to represent speaker timbre. And in variable-length speaker modeling [25, 24, 23], scale-dot attention and cross-attention are usually utilized to fuse speaker and content representations. Previous studies [21, 16, 20] have explored the correlation between channel and speaker timbre information in zero-shot VC, but only at the utterance-level and single channel dimension. In contrast, our proposed method focuses on fine-grained speaker modeling and performs speaker retrieval in both temporal and channel dimensions.
@@ III-A The Framework of MTCR-VC
In this paper, we introduce multi-level temporal channel retrieval based voice conversion (MTCR-VC), a novel approach that integrates time-varying and multi-level speaker modeling frameworks with substantial improvements, as depicted in Figure 2. Based on the structure of FragmentVC [23], MTCR-VC disentangles speech into speaking style, linguistic content, and speaker timbre using a speaking style module, a content encoder, and an MTCR speaker module, respectively. The speaking style and linguistic content representations are learned from the pitch contour $P_{src}$ and BNF $B_{src}$ of the source speech. Then the content, pitch, and rhythm representations are added to form $Z_{add}$. The MTCR speaker module extracts multiple fine-grained speaker representations from the target speaker’s speech. Finally, the speech decoder progressively combines the learned speech representations to produce the corresponding target Mel spectrogram $M_{trg}$, which conveys the target speaker’s timbre while maintaining the speaking style and linguistic content of the source speech.
@@ III-A1 Content Encoder
The content encoder is a Conformer-like structure that learns the linguistic content representation, as indicated in the orange section on the left of Fig. 2. Here the bottleneck features (BNF) serves as the input for the content encoder, as BNF is considered as a speaker-irrelevant linguistic representation [3, 34], obtained from the encoder of a well-trained ASR model.
@@ III-A2 Speaking Style Module
Based on the above structure, we propose the MTCR speaker module and a training strategy, which are elaborated in Section III-B and III-D. Moreover, we make modifications to the speech decoder [23] to facilitate the integration of the MTCR speaker module, which are also detailed in Section III-C.
@@ III-B MTCR Speaker Module: Speaker Modeling
Speaker timbre is a dynamic characteristic that varies across temporal and channel regions of speech at different granularities. Inspired by the representation learning ability of attention mechanism [30, 42], we propose the multi-level temporal-channel retrieval (MTCR) speaker module for retrieving speaker-related information from the target speaker’s speech in both temporal and channel regions with different granularities. Specifically, as shown in the blue part of Fig. 2, the MTCR speaker module comprises multiple TCR blocks to achieve multi-granularity speaker retrieval. Each TCR block sequentially performs temporal and channel retrieval at a specific granularity. And the speaker embedding (x-vector) derived from a SV model [32] is considered an ideal query to guide the speaker retrieval since the SV model achieves high accuracy in speaker verification and learns accurate speaker discriminative ability.
@@
Temporal Retrieval. The objective of our work is to enable the speaker module to know when it should pay more attention to capture speaker information. For this purpose, we adopt temporal retrieval. As speech varies in duration and speaker timbre exhibits different characteristics at various temporal granularities, temporal segmentation, in which the feature are equally divided into several groups along the temporal dimension, has to considered before the application of attention mechanism. Specifically, after the convolution layer processing the input $Z_{s_{l-1}}$ into the output $H_{t_{l}}$, we perform temporal segmentation on $H_{t_{l}}$ using a temporal scale factor $\gamma_{t}$. Then the resulting segmentation $H^{{}^{\prime}}_{t_{l}}\in\mathbb{R}^{\frac{T}{\gamma_{t}}\times\gamma_{t}\times C}$ is used as the key $K_{t_{l}}$ and value $V_{t_{l}}$. Meanwhile, the x-vector $S_{spk}$ is transformed by an FC layer and used as the query $Q_{t_{l}}\in\mathbb{R}^{1\times C}$. By measuring the similarity between $Q_{t_{l}}$ and $K_{t_{l}}$, we obtain an attention map $A_{t_{l}}\in\mathbb{R}^{\frac{T}{\gamma_{t}}\times 1\times\gamma_{t}}$ that captures the distribution of speaker timbre in each temporal segment. Finally, we multiply $A_{t_{l}}$ and $V_{t_{l}}$, yielding $H_{c_{l}}\in\mathbb{R}^{\frac{T}{\gamma_{t}}\times C}$ that we also use for the following channel retrieval. The entire process of temporal retrieval can be summarized as follows:
@@
Channel Retrieval. To retrieve speaker timbre from speech, it is crucial to not only determine when the speaker timbre is present but also where it exists in the channel dimension. Therefore, we use both temporal and channel retrieval techniques to achieve this. Fig. 3 illustrates our approach, which begins with channel segmentation applied to $H_{c_{l}}$. Since the speaker timbre changes with time-varying style or content, the distribution of speaker timbre on the channel is also related to the temporal changes. Therefore, we consider the temporal range $\gamma_{tr}$ of each channel segment during the channel segmentation. After the channel segmentation, the result $H^{{}^{\prime}}_{c_{l}}$ is in the shape of $\{\frac{T^{{}^{\prime}}}{\gamma_{tr}},\frac{C}{\gamma_{c}},\gamma_{c},\gamma_{tr}\}$, where $T^{{}^{\prime}}$ represents $\frac{T}{\gamma_{t}}$ mentioned above, and $\gamma_{c}$ represents the channel scale factor. Then, $H^{{}^{\prime}}_{c_{l}}$ is used as the key $K_{c_{l}}$ and value $V_{c_{l}}$. Meanwhile, after processing by an FC layer, the x-vector $S_{spk}$ is used as the query $Q_{c_{l}}$. To compute the attention map $A_{c_{l}}\in\mathbb{R}^{\frac{T^{{}^{\prime}}}{\gamma_{tr}}\times\frac{C}{\gamma_{c}}\times 1\times\gamma_{c}}$ for each channel segment, we measure the similarity between $K_{c_{l}}$ and $Q_{c_{l}}$. Finally, we obtain the TCR block result $Z_{s_{l}}$, with dimensions of $\{\frac{T}{\gamma_{t}},\frac{C}{\gamma_{c}}\}$, which is computed by multiplying $A_{c_{l}}$ and $V_{c_{l}}$.

### v3

@@ Abstract
Zero-shot voice conversion (VC) converts source speech into the voice of any desired speaker using only one utterance of the speaker without requiring additional model updates. Typical methods use a speaker representation from a pre-trained speaker verification (SV) model or learn speaker representation during VC training to achieve zero-shot VC. However, existing speaker modeling methods overlook the variation of speaker information richness in temporal and frequency channel dimensions of speech. This insufficient speaker modeling hampers the ability of the VC model to accurately represent unseen speakers who are not in the training dataset. In this study, we present a robust zero-shot VC model with multi-level temporal-channel retrieval, referred to as MTCR-VC. Specifically, to flexibly adapt to the dynamic-variant speaker characteristic in the temporal and channel axis of the speech, we propose a novel fine-grained speaker modeling method, called temporal-channel retrieval (TCR), to find out when and where speaker information appears in speech. It retrieves variable-length speaker representation from both temporal and channel dimensions under the guidance of a pre-trained SV model. Besides, inspired by the hierarchical process of human speech production, the MTCR speaker module stacks several TCR blocks to extract speaker representations from multi-granularity levels. Furthermore, we introduce a cycle-based training strategy to simulate zero-shot inference recurrently to achieve better speech disentanglement and reconstruction. To drive this process, we adopt perceptual constraints on three aspects: content, style, and speaker. Experiments demonstrate that MTCR-VC is superior to the previous zero-shot VC methods in modeling speaker timbre while maintaining good speech naturalness.
@@ I Introduction
Voice conversion (VC) aims to convert the source speech to a target speaker without changing the linguistic content. VC has been deployed to many applications, including dubbing, live broadcast, voice anonymization, and pronunciation correction. In recent years, neural networks-based VC systems, such as generative adversarial network (GAN) [1], variational auto-encoder (VAE) [2], and recognition-synthesis autoencoder framework [3] are widely adopted. However, these VC models can only convert source speech to that of predefined target speakers. Meanwhile, these VC models usually require many speech recordings of the target speaker for model training. As data collection is expensive and time-consuming, building a high-quality VC system with minimal data requirements is more practical for real-world applications. Consequently, zero-shot VC, also called any-to-any VC, which aims to convert source speech to that of any speaker given only one utterance of the speaker, has drawn much attention recently. In this paper, we focus on the problem of zero-shot VC.
@@
As only one utterance is available, one of the key challenges in zero-shot VC is effectively capturing the target speaker’s timbre to achieve good speaker similarity in the converted speech. Using a look-up table (LuT) to represent speaker identity [4, 5, 6] is a common practice but is limited to predefined speakers in the training data. Many recent studies attempt to directly capture speaker timbre from the limited speech of the target speaker. An intuitive approach is to leverage a pre-trained robust speaker verification (SV) model to extract utterance-level speaker representation [7, 8, 9, 10]. However, the SV model trained on many speakers and recording conditions is optimized for speaker classification but not for the perceptual speaker similarity in the sense of human hearing. Instead of using an external SV model, many studies [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21] decompose speech into linguistic content and speaker timbre, and even speaking style, aiming to separate utterance-level speaker timbre from other speech components. In the above approaches, speaker timbre is considered static and time-independent and modeled as a single coarse-grained fixed-length vector. To capture the dynamically varying speaker characteristics within an utterance, some recent studies [22, 23, 24, 25, 26] focus on modeling more fine-grained speaker representation, which extracts speaker timbre from multiple aspects, covering multi-level and time-varying speaker representation. Based on the U-net [27] structure, Li et al. [28], Wu et al. [22], and Li et al. [26] extract utterance-level speaker representations from multiple stacking layers and feed them to corresponding decoder layers. To access time-varying speaker information, some studies [25, 23, 24] extract variable-length speaker representation and fuse it into the converted speech according to the content-based alignment between source speech and target speaker speech.
@@
While the above progress has been made, existing methods of modeling fine-grained target speaker timbre are insufficient to capture the dynamic variation characteristic of speaker timbre. The above speaker representation is usually extracted from the time-frequency space of speech, where a speech spectrogram includes both temporal and frequency channel dimensions. There are differences in speaker timbre richness in both temporal and channel dimensions of speech. Insufficient speaker modeling in these two dimensions may cause the unstable performance of the VC model to represent unseen speakers. Moreover, speech production research [29] shows that the frequency distributions of speech from different speakers lead to varying speaker timbre information in frequency channels [30, 31, 32]. Speech content, such as vowels, consonants, and para-linguistic features, carry distinct speaker timbre information reflected in temporal and frequency channel dimensions while silent speech segments apparently convey no speaker timbre information [25]. On the other hand, the human speech production mechanism is hierarchical [33, 29] in nature, from long-term airflow generation to fine-grained phoneme-related articulator movements and vocal filtering. Speaker-related information thus varies at different stages of speech production with different temporal and channel granularities. Finally, different speech factors such as linguistic content, speaking style, and speaker timbre are highly entangled in speech, and factor disentanglement is necessary for better speaker timbre extraction. Compared with utterance-level speaker modeling, time-varying fine-grained speaker modeling can transfer more speaker information from the target speaker’s speech to the converted speech. But this exacerbates the difficulty of disentanglement [25, 23].
@@
To address the aforementioned problems in zero-shot VC, we propose a novel fine-grained speaker modeling method called temporal-channel retrieval (TCR), which captures the dynamic variation of speaker timbre in both temporal and channel dimensions. Specifically, with the help of the attention mechanism, the speaker embedding from the pre-trained SV model is used as a query to retrieve speaker timbre information in variable-length speaker representation from the target speaker. Inspired by the hierarchical nature of speech production, we employ multi-level TCR (MTCR) in an encoder-decoder based U-net structure for voice conversion [27] where fine-grained speaker information in different granularities is retrieved from different encoder layers. In the decoder, these multi-level speaker representations are then fused with content and style representations from the source speech to generate target speech. Furthermore, to perform better speech disentanglement, we introduce a cycle-based training strategy to simulate zero-shot inference in a recurrent fashion. perceptual constraints [34] on three aspects, including content, style, and speaker, are adopted to drive this process. Zero-shot voice conversion experiments show that the proposed MTCR-VC approach generalizes well to cross-set speakers with superior speaker similarity. Audio samples can be found on our demo page. [^1] https://kerwinchao.github.io/demo_zslvc/
@@
  • We propose a novel multi-level framework, MTCR-VC, for fine-grained speaker modeling in zero-shot VC. The multi-level temporal-channel retrieval is designed to extract fine-grained speaker representations under the query of SV’s speaker embedding in both temporal and channel dimensions and hierarchically capture speaker timbre from different granularities.
@@ II Related work
This section will review related works on voice conversion and the typical zero-shot VC frameworks. We will also review the utilization of the attention mechanism in VC.
@@ II-A Voice Conversion
VC approaches can be categorized into one-to-one, many-to-many, or any-to-any VC based on the number of speakers that the VC model supports. Conventional VC approaches focus on one-to-one VC, which align acoustic features of parallel data between a pair of source-target speakers and perform frame-wise mapping using Gaussian mixture models [35, 36] and neural networks [37, 38]. With the high cost of parallel data, recent studies mainly use non-parallel data for voice conversion. GAN [4, 5] and VAE [1] are further proposed to support many-to-many VC. With the help of text supervision, phonetic posteriorgram (PPG) and neural bottleneck feature (BNF) [3] computed from an automatic speech recognition (ASR) model are assumed to be speaker-independent content representations. And thus PPG- and BN-based recognition-synthesis framework is widely used in any-to-many VC. Besides, many disentanglement methods based on information bottleneck [7], perturbation [8, 16], and representation constraints [17, 13] decompose speech into content, speaker timbre, and speaking style to achieve any-to-any VC, also called zero-shot VC. Based on the recognition-synthesis framework, this work proposes a new method for zero-shot VC.
@@ II-B Typical Zero-shot VC Framework
The key to voice conversion is to decompose speech into linguistic content and speaker timbre or even speaking style components. Ideally, only the content and style components from the source speech are transferred to the target speech with the target speaker’s timbre. In many studies [11, 12, 23, 10], style is not explicitly considered in the disentanglement. In zero-shot scenarios, the speaker timbre representation is extracted from a short speech utterance of the target speaker, and then the speaker representation is fused with the extracted content and style of the source speech to generate the target speech. The extraction of speaker timbre can be done at utterance level or fine-grained level, as shown in Fig. 1.
@@ II-B2 Fine-grained Speaker Modeling
Fine-grained speaker modeling differs from utterance-level speaker modeling in that it accounts for changes in speaker timbre over time and considers dynamic variation characteristics. Fig. 1(b) and (c) illustrate two categories of fine-grained speaker modeling: time-varying and multi-level. In the time-varying speaker modeling [25, 23, 24] as shown in Fig.1(b), variable-length speaker representation is extracted from the target speaker’s speech and then fused with the content and style of the source speech based on the alignment between them. On the other hand, multi-level speaker modeling [28, 26] shown in Fig.1(c) typically adopts a U-net structure to extract multiple utterance-level speaker representations from stacked layers and then feeds them to the corresponding decoder layers. Recent studies [23, 25, 24] attempt to simultaneously employ time-varying and multi-level speaker modeling.
Despite this progress, existing methods are insufficient to capture the dynamic variation characteristic of the speaker timbre, which varies across temporal and channel regions of speech at different granularities. This insufficient may cause unstable speaker modeling and result in low speaker similarity. Besides, in a time-varying fine-grained speaker modeling framework, speech factors are much easier to entangle with each other, making achieving zero-shot more challenging. This paper focuses on fine-grained speaker modeling in our proposed zero-shot VC framework, which models speaker timbre from both temporal and channel dimensions at different granularities. Comprehensive representations of speech components and a well-designed training strategy ensure the decoupling capabilities of our model.
@@ II-C Attention Mechanism in Speaker Modeling
In recent years, many attention mechanisms have been proposed, such as spatial attention [39], channel attention [40], temporal attention [41], and their combination [42, 43]. Some attention mechanisms have also been explored in speaker modeling. In Desplanques et al. [32], Squeeze-and-excitation [40] block (SE) is introduced to aggregate speaker information along the channel axis. Based on Desplanques et al. [32], Liu et al. [30] introduce multi-scale channel attention to better capture the speaker characteristics at any local frequency region. Besides, Sang et al.[31] improve the process of extracting utterance-level speaker representation from frame-level speaker representation by modified channel attention driven by a discrete cosine transform. For zero-shot VC, following Desplanques et al. [32], Du et al. [21], and Choi et al. [16] utilize SE block to improve the speaker modeling ability for zero-shot VC. CA-VC uses channel attention to improve content learning and encourage utterance-level speaker modeling. Besides, Retriever [24] uses cross-attention [44] to learn a set of permutation invariant tokens to represent speaker timbre. In variable-length speaker modeling [25, 24, 23], scale-dot attention and cross-attention are usually utilized to fuse speaker and content representations. Previous studies [21, 16, 20] have explored the correlation between channel and speaker timbre information in zero-shot VC, but only at the utterance level and single channel dimension. In contrast, our proposed method focuses on fine-grained speaker modeling and performs speaker retrieval in both temporal and channel dimensions.
@@ III-A The Framework of MTCR-VC
In this paper, we introduce multi-level temporal channel retrieval-based voice conversion (MTCR-VC), a novel approach that integrates time-varying and multi-level speaker modeling frameworks with substantial improvements, as depicted in Figure 2. Based on the structure of FragmentVC [23], MTCR-VC disentangles speech into speaking style, linguistic content, and speaker timbre using a speaking style module, a content encoder, and an MTCR speaker module, respectively. The speaking style and linguistic content representations are learned from the source speech’s pitch contour $P_{src}$ and BNF $B_{src}$. Then the content, pitch, and rhythm representations are added to form $Z_{add}$. The MTCR speaker module extracts multiple fine-grained speaker representations from the target speaker’s speech. Finally, the speech decoder progressively combines the learned speech representations to produce the corresponding target mel spectrogram $M_{trg}$, which conveys the target speaker’s timbre while maintaining the source speech’s speaking style and linguistic content.
@@ III-A1 Content Encoder
The content encoder is a Conformer-like structure that learns the linguistic content representation, as indicated in the orange section on the left of Fig. 2. Here, the bottleneck features (BNF) serve as the input for the content encoder, as BNF is considered as a speaker-irrelevant linguistic representation [3, 34], obtained from the encoder of a well-trained ASR model.
@@ III-A2 Speaking Style Module
Based on the above structure, we propose the MTCR speaker module and a training strategy, which are elaborated in Section III-B and III-D. Moreover, we modify the speech decoder [23] to facilitate the integration of the MTCR speaker module, which are also detailed in Section III-C.
@@ III-B MTCR Speaker Module: Speaker Modeling
Speaker timbre is a dynamic characteristic that varies across temporal and channel regions of speech at different granularities. Inspired by the representation learning ability of attention mechanism [30, 42], we propose the multi-level temporal-channel retrieval (MTCR) speaker module for retrieving speaker-related information from the target speaker’s speech in both temporal and channel regions with different granularities. Specifically, as shown in the blue part of Fig. 2, the MTCR speaker module comprises multiple TCR blocks to achieve multi-granularity speaker retrieval. Each TCR block sequentially performs temporal and channel retrieval at a specific granularity. The speaker embedding (x-vector) derived from an SV model [32] is considered an ideal query to guide the speaker retrieval since the SV model achieves high accuracy in speaker verification and learns accurate speaker discriminative ability.
@@
Temporal Retrieval. The objective of our work is to enable the speaker module to know when it should pay more attention to capture speaker information. For this purpose, we adopt temporal retrieval. As speech varies in duration and speaker timbre exhibits different characteristics at various temporal granularities, temporal segmentation, in which the features are equally divided into several groups along the temporal dimension, must be considered before applying the attention mechanism. Specifically, after the convolution layer processing the input $Z_{s_{l-1}}$ into the output $H_{t_{l}}$, we perform temporal segmentation on $H_{t_{l}}$ using a temporal scale factor $\gamma_{t}$. Then the resulting segmentation $H^{{}^{\prime}}_{t_{l}}\in\mathbb{R}^{\frac{T}{\gamma_{t}}\times\gamma_{t}\times C}$ is used as the key $K_{t_{l}}$ and value $V_{t_{l}}$. Meanwhile, the x-vector $S_{spk}$ is transformed by an FC layer and used as the query $Q_{t_{l}}\in\mathbb{R}^{1\times C}$. By measuring the similarity between $Q_{t_{l}}$ and $K_{t_{l}}$, we obtain an attention map $A_{t_{l}}\in\mathbb{R}^{\frac{T}{\gamma_{t}}\times 1\times\gamma_{t}}$ that captures the distribution of speaker timbre in each temporal segment. Finally, we multiply $A_{t_{l}}$ and $V_{t_{l}}$, yielding $H_{c_{l}}\in\mathbb{R}^{\frac{T}{\gamma_{t}}\times C}$ that we also use for the following channel retrieval. The entire process of temporal retrieval can be summarized as follows:
@@
Channel Retrieval. To retrieve speaker timbre from speech, it is crucial to not only determine when the speaker timbre is present but also where it exists in the channel dimension. Therefore, we use both temporal and channel retrieval techniques to achieve this. Fig. 3 illustrates our approach, which begins with channel segmentation applied to $H_{c_{l}}$. Since the speaker timbre changes with time-varying style or content, the distribution of speaker timbre on the channel is also related to the temporal changes. Therefore, we consider each channel segment’s temporal range $\gamma_{tr}$ during the channel segmentation. After the channel segmentation, the result $H^{{}^{\prime}}_{c_{l}}$ is in the shape of $\{\frac{T^{{}^{\prime}}}{\gamma_{tr}},\frac{C}{\gamma_{c}},\gamma_{c},\gamma_{tr}\}$, where $T^{{}^{\prime}}$ represents $\frac{T}{\gamma_{t}}$ mentioned above, and $\gamma_{c}$ represents the channel scale factor. Then, $H^{{}^{\prime}}_{c_{l}}$ is used as the key $K_{c_{l}}$ and value $V_{c_{l}}$. Meanwhile, after processing by an FC layer, the x-vector $S_{spk}$ is used as the query $Q_{c_{l}}$. To compute the attention map $A_{c_{l}}\in\mathbb{R}^{\frac{T^{{}^{\prime}}}{\gamma_{tr}}\times\frac{C}{\gamma_{c}}\times 1\times\gamma_{c}}$ for each channel segment, we measure the similarity between $K_{c_{l}}$ and $Q_{c_{l}}$. Finally, we obtain the TCR block result $Z_{s_{l}}$, with dimensions of $\{\frac{T}{\gamma_{t}},\frac{C}{\gamma_{c}}\}$, which is computed by multiplying $A_{c_{l}}$ and $V_{c_{l}}$.